hundreds of small cube satellites called CubeSats are being launched experimentally every year. NASA limits the altitude of these shoe box sized satellites to 700km so that the air friction will eventually slow them down and cause them to fall to earth within 25 years. what is the force of gravitational attraction between one of these 1.33kg satellites and the earth? the mass of the earth is 5.98 x 1024kg and the radius of the earth is 6.4 x 106 meters


1
Expert's answer
2021-03-26T11:38:09-0400
"F=G\frac{mM}{r^2}\\ F=(6.67\cdot10^{-11})\frac{(5.98\cdot10^{24})(1.33)}{(6.4\cdot10^{6}+7\cdot10^{5})^2}=10.5\ N"